# Inga Strümke5, Vince I. Madai6,7, Tobias Budig8, Ali Sunyaev9, and Adam Hilbert10 On behalf of the VALIDATE consortium

### v1

@@ Abstract
The joint implementation of Federated learning (FL) and Explainable artificial intelligence (XAI) will allow training models from distributed data and explaining their inner workings while preserving important aspects of privacy. Towards establishing the benefits and tensions associated with their interplay, this scoping review maps those publications that jointly deal with FL and XAI, focusing on publications where an interplay between FL and model interpretability or post-hoc explanations was found.
In total, 37 studies met our criteria, with more papers focusing on explanation methods (mainly feature relevance) than on interpretability (mainly algorithmic transparency). Most works used simulated horizontal FL setups involving 10 or fewer data centers. Only one study explicitly and quantitatively analyzed the influence of FL on model explanations, revealing a significant research gap. Aggregation of interpretability metrics across FL nodes created generalized global insights at the expense of node-specific patterns being diluted. 8 papers addressed the benefits of incorporating explanation methods as a component of the FL algorithm.
Studies using established FL libraries or following reporting guidelines are a minority. More quantitative research and structured, transparent practices are needed to fully understand their mutual impact and under which conditions it happens.
@@
Machine Learning, Artificial Intelligence (AI), Explainable AI, Model Interpretability, Federated Learning, Data Privacy Preservation.
@@ I Introduction
The development of trustworthy AI systems depends on multiple different ethical principles like privacy preservation and explicability [1, 2]. Ensuring that AI systems address such principles is crucial for maintaining trust, compliance, and ethical standards. Such principles are critical in data-sensitive applications such as banking [3] and healthcare [4, 5].
Privacy concerns are the main reason why healthcare data for AI models, such as patient outcome prognosis, mainly originate from single institutions. Anonymizing extensive healthcare data effectively is challenging due to the high risk of re-identifying patients [6]. Allowing models to be trained across multiple institutions without sharing raw data, for example, with federated learning (FL), addresses this issue. FL has gained prominence as a solution to train models from distributed data while preserving these privacy aspects [7, 8]. Specifically, FL enables the training of machine learning (ML) models from distributed data that are available to different users or institutions, without data being explicitly transferred. Therefore, FL solves certain privacy-related governance issues as it restricts access to sensitive information from each institution while still gaining insights from all available data [9, 10]. Using data from different institutions can lead to more diverse training data, which reduces the bias of ML models and improves generalizability compared to models trained on homogeneous data [11]. The three main categories within FL are horizontal FL (HFL), vertical FL (VFL), and federated transfer learning (FTL). HFL involves data scattered across the sample space, where multiple parties have different subsets of data points but all data have the same feature space. VFL allows learning from data distributed across different different feature spaces for the same data sample. FTL focuses on transferring knowledge across different federated learning settings [12].
Another ethical principle often advocated by ethicists and policymakers in high-risk AI domains such as healthcare is explicability. This means that: a) processes must be transparent, b) the capabilities and purpose of AI systems must be openly communicated, and c) one must be able to explain the decisions and predictions made by AI systems to those directly or indirectly affected by them [1, 2]. When a model’s inference or decision-making process is transparent and understandable, it is more likely to be trusted and accepted by stakeholders, such as customers, regulators, and users [13].
As stated by the European Union’s regulations, aspects (a) and (c) of explicability, whose technical manifestation is known as explainability [14], need to be provided through additional information — next to the performance of the AI system — on how an AI system arrives at a certain output [15]. The concept of explainable artificial intelligence (XAI) aims to explain the inner workings of ML algorithms and increasing model comprehension. XAI is thereby concerned with the development of algorithms and methods that achieve explainability [16]. Explicability can be achieved by either training interpretable ML models, or by using post-hoc explanation methods (i.e., (a) and (c) respectively regarding the European Union guideline).
Unfortunately, XAI literature has not agreed on one definition of explainability. The terms “explainability”, “explicability”, and “interpretability” are often used interchangeably [17]. In this work, we adopt the definitions provided by [15], which include interpretable modeling and explanation methods, as detailed in the following paragraphs.
An ML model is said to be interpretable if its decision process can be inherently and intuitively understood by the intended user [15]. The word ‘inherently‘ is crucial in this definition, as interpretability of a given model is defined as a passive property that reflects the level at which such a model makes sense for a human observer [18] without the need to perform extra computation. Examples of interpretable ML models are linear and logistic regression models, where variable importance can be inferred by investigating model weights, or decision trees, which can be intuitively followed by humans [15].
In contrast, post-hoc explanation methods try to explain non-interpretable models actively by performing additional computations with the intent of clarifying or detailing their internal functions [18, 15]. Such computations include evaluating or calculating gradients of the original model for a set of differently perturbed inputs [19]. With that, AI systems can calculate the importance of features [20], the dependence on feature variations of the output [21], or the minimal input change that causes a label flip [22]. An explanation method, therefore, is a comprehensible interface between humans and a decision maker while simultaneously providing approximations of the decision-maker [23].
Extensive research has been conducted on FL and XAI separately. Despite the importance of their combined use for creating trustworthy AI systems, it remains unclear whether these two technologies complement each other, present conflicting requirements, or can be addressed separately. FL and XAI coexist in many works, but few works discuss the impact of one on the other. Furthermore, the existing literature inadequately addresses the methodologies, challenges, and outcomes related to the integration of FL and XAI, leaving a critical gap in understanding their mutual influence.
On one hand, XAI can mitigate the risks associated with training ML systems in federated settings. Learning from heterogeneous data can inadvertently lead to spurious correlations learned by the ML model, as the decentralized data might not represent the overall distribution and contain biases specific to subsets of the data. Explainability can help practitioners determine whether the model’s predictions are based on valid patterns rather than misleading correlations, not propagating undesired biases and, therefore, adhering to ethical standards. Additionally, explainability can help identify malicious agents trying to poison the FL process.
On the other hand, the potential risks associated with the interplay between these technologies must be documented. There is no unified effort to determine, for example, whether FL leads to reduced interpretability or less accurate explanations. Also, explanations could make the FL network more vulnerable to attacks. Moreover, preliminary research [24] indicates differences in model explanations between FL and traditional centralized models, suggesting unique challenges in applying XAI in FL contexts. This underscores the need for further research on how FL influences the contribution of input variables and the effectiveness of XAI methods.
@@
To understand the interplay between FL and XAI, it is crucial to consider the findings, setups, and conditions under which these studies have been conducted. This scoping review aims to map both experimental and methodological studies that examine the interaction between FL and XAI. Since FL can potentially influence both model interpretability and post-hoc explanations, this review paper investigates both concepts of XAI. It focuses on research articles that either propose FL methods incorporating explainability, discuss practical applications where FL and XAI are jointly used, or analyze the interplay between these technologies through empirical studies. Given the fragmented and emerging nature of research at this intersection, this review provides a comprehensive overview of current work, identifies key concepts and evidence types, and highlights gaps to guide future research efforts.
@@
The remainder of this paper is structured as follows. Sec. II introduces the related work. Sec. III describes the methodology we followed. A more detailed analysis of the interplay between FL and XAI found in the reviewed papers is given in Sec. IV. A quantitative analysis of the extracted information is shown in Sec. V. The results are discussed in Sec. VI, and Sec. VII concludes the paper. The Appendix contains summarized information about how all included papers jointly with FL and XAI.
@@ II Related Work
Review papers on the fields of XAI and FL separately are very common, and there are even meta-reviews [25] and systematic reviews from different points of view (see, e.g. [26, 27] regarding FL applications for biomedical data). An extensive review on XAI methods can be found in [16] or [28]. Even though, there are many fewer reviews concerning the joint application of XAI and FL, none analyzes the interplay in depth, which motivates our work.
@@ II-A Reviews about FL and XAI
The review in [29] introduces Fed-XAI, which involves learning interpretable models in federation, in addition to explanation methods applied on any federated trained model. They used [18] as a source for XAI taxonomy and highlight a general diversity in definitions used in the studied literature. Additionally, they gave an overview of the “current status in Fed-XAI” by discussing a set of works relating to the two concepts with no analysis on the interplay between FL and XAI.
The review in [30] gathers relevant works concerning the “FED-XAI” concept, defined by them as a discipline “which aims to bring together these two approaches into one”. Differently from our search terms, which include variations of the words “explanation” and “interpretability”, their search is limited to the explicit string “FED-XAI”, returning a reduced number of publications. The authors do not provide a comparative analysis of the interplay between FL and XAI among the found papers.
The paper [31] surveys “interpretable federated learning” and propose an interpretable FL taxonomy that enables learning models to explain prediction results, support model debugging, and provide insights into data owner contributions. The survey analyzes representative interpretable FL approaches, commonly adopted performance evaluation metrics, and future research directions. The main focus is on how to use interpretable methods to improve the FL process, but not on the impact of FL on the explanations of the predictions.
In contrast to these works, our review 1) applies a scoping review methodology, 2) provides a more detailed set of results, 3) summarizes how the reviewed works jointly deal with FL and XAI, and 4) analyzes the findings on the interplay between them.
II-B Contribution-aware FL
A small but salient subset of the approaches investigating the intersection of FL and XAI has utilized feature relevance methods (e.g. Shapley values) to measure the contribution of each client participating in a FL process. These approaches, commonly referred to as contribution-aware FL, aim to incentivize participants to engage during model training and to fairly distribute rewards based on contributions. One of the first approaches to contribution-aware FL suggested using Shapley values to interpret contributions in FL networks in [32]. The contributions are diverse, ranging from node liability [33] to biases within data sets [34]. Recent approaches have also extended SHAP-based [20] contribution determination to provide visualizations to evaluate data privacy [35] and to improve the reliability of prediction via clustering patients [36]. Such methods do not necessarily constitute an influence of XAI on the FL process and are therefore not central to the present review.
@@ III Methods
The present scoping review addresses the concept of joint application of FL and XAI and their potential interplay. Two research types have been addressed, namely the methodological advances and the experimental results.
We aimed at examining the range and nature of research activity on the topic, summarizing the research findings, and identifying research gaps, following a standard scoping review methodology [37]. The presence of two subconcepts of explainability (interpretable models and post-hoc explanation), the two research types (methodological and experimental), and the diversity of joint approaches to FL and XAI increased the complexity of the study. Understanding and summarizing the diverse approaches required a deeper analysis of each included paper. A more detailed description of the methodology is available in the pre-published research protocol [38].
In our survey, the time frame of the published papers was between 2019 and April 2023 (date of protocol pre-publication). No relevant papers were found before 2019.
III-A Research questions
The research questions of this work address both the training of interpretable ML models using FL, and the explanation methods applied to ML models trained via FL. Furthermore, the questions were categorized into two groups:
@@ III-A1 Methodological advances
  • Has any work proposed an FL method that takes into account the ulterior application of explanation methods?
@@ III-A2 Experimental results
  • In which contexts (fields of application) have results been obtained regarding FL and XAI?
@@
  • Has any work compared the effects of FL on the interpretability of the resulting models against a centralized learning algorithm?
  • Has any work compared the explanations obtained from a model trained via FL against a centralized learning algorithm?
We used the following search engines: Google Scholar, IEEExplore, PubMed, Scopus, and Web of Science. The terminology issue described in the introduction influenced the search strategy, which aimed at identifying all papers mentioning FL jointly with a word related to either interpretability or explainability. The search terms are detailed in Table I.
@@
IEEEXplore (“federated”) AND (learning) AND ((explainable) OR (interpretable) OR (explaining) OR (explainability) OR (interpreting) OR (interpretability) OR (interpret))
@@
PubMED (“federated”) AND (learning) AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)
Scopus TITLE-ABS-KEY(“federated” AND learning AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)) AND (LIMIT-TO(DOCTYPE, “ar”) OR LIMIT-TO(DOCTYPE, “cp”))
@@
III-B Inclusion/exclusion criteria
We initially identified all research articles published in peer-reviewed conferences and journals and pre-published works made available in preprint services such as arXiv over the past three years. We required that the full text was available.
To determine their inclusion in the review, papers had to either discuss in depth the relation and interaction between XAI and FL, or apply both technologies jointly in a practical setting. To rule out papers that mentioned both technologies without detailed discussion or practical application, we performed a screening procedure, in which the article must answer positively to at least one of the following screening questions:
  • SQ1) Does the paper propose a novel method integrating FL and XAI?
  • SQ2) Does the paper report results from experiments applying FL and XAI in a real dataset?
  • SQ3) Does the paper discuss or assess the impact of FL on explanations or model interpretability?
The exclusion criteria were defined as follows. Theses were excluded. Reviews/meta-reviews/surveys were excluded. Preprints published more than 3 years before the protocol prepublication, for which a final published version did not exist, were excluded.
III-C Extraction
To facilitate efficient and parallel data extraction, we utilized a shared online spreadsheet. All authors were involved in the data extraction which allowed us double-coding of all information. The extracted items were detailed in the protocol [38], and a summarized listed is provided here:
@@ III-C Extraction
  • Nomenclature: whether the paper provides or cites a definition of explainability, and whether the nomenclature regarding explainability coincides with the one introduced here.
  • Nature of the paper: whether it is general/theoretical or applied; if applied, what field it is applied on, and whether the proposed idea is practically validated in the field of application.
  • Data characteristics: information modality, type and amount (number of unique data samples) of the data used in the experiments.
@@
  • XAI-specific characteristics: whether an interpretable model or an explanation method is used [^1] If the nomenclature in a papers does not agree with the one mentioned in the Introduction, we identify whether it uses explanation methods or interpretable models according to our understanding., its type, and the specific explanation method (e.g.: SHAP, GradCAM [39]) or way of interpreting the model (e.g., weights at the first layer of DNN).
  • Interplay between FL and XAI: a paragraph summarizing how the paper deals jointly with FL and XAI; influence of FL on explanations or interpretability (e.g. significant differences in variable importance ratings between centrally learned and model trained via FL); influence of XAI onto federated training (e.g., modified merging step in central server after collecting locally trained models); and whether this influence is quantified, and how.
@@
During both the screening and the extraction process, double coding was practiced. Each screened paper was independently assessed by two different authors, and answers to the extraction points for each of the included papers were obtained by two different authors, resolving differences by discussion between the two involved authors; disagreements were solved by discussion within the entire author consortium. While extracting, we distinguished between explanation methods and model interpretability according to the nomenclature above introduced, even though some papers did not agree with the latter.
A flowchart summarizing the number of papers found, excluded, and included in this study, along the lines of [40], is displayed in Fig. 2.
IV Interplay between FL and XAI
This section explores studies that have either integrated explainability into FL methods or analyzed how one technology influences the other. Studies that jointly deal with FL an XAI but do not necessarily report an influence are discussed in the Appendix. Here, we first discuss instances where FL has impacted post-hoc explanations (Sec. IV-A). Next, we examine works where explanation methods have influenced FL training as a design step in the FL process (Sec. IV-B). Lastly, we cover studies where FL has altered model interpretability, or the implementation of interpretable models has influenced FL training (Sec. IV-C).
IV-A FL impacting XAI
Among all papers reviewed, only [41] quantitatively analyzes the impact of FL on explanations. Additionally, [42] highlights privacy concerns associated with an explanation method used on FL models.
An FL procedure is developed in [41] for taxi travel-time prediction, based on time-series and geographical data. The authors develop a federated feature attribution aggregation method and test the similarity of FL-calculated explanations compared to those calculated centrally. Several XAI techniques are tested, specifically DL-specific attribution explanation methods such as DeepLIFT [43], Saliency [44], Input X Gradient [45], Guided Backpropagation [46], Deconvolution [47], and Layer-wise Relevance Propagation [48]. It is reported that, upon training via FL, explanations across different local models lead to different variable importance rankings. However, no significant difference was found in federated feature attributions in FL compared to those in centrally trained models.
An FL model is trained in [42] to predict the latency in the creation of a network slice in a mobile communications scenario. Each slice manager provides data regarding CPU/RAM capacity and usage, serving as an FL node. The models are evaluated both locally and globally using SHAP [20], LIME [49], Partial Dependence Plots (PDP) [50, Ch. 8.1], and RuleFit [51]. For the PDP technique, the FL clients only share plots that display the percentage of feature impact on the target label to preserve data privacy. The influence is analyzed qualitatively with a focus on privacy concerns. Overall, it is shown that PDP explanations raise privacy concerns since they are executed on the client side.
IV-B Explanation method impacts FL training
Among the papers that reported that the use of an explanation method impacts the FL training, a salient subset uses XAI to defend the FL process from the negative effects of defaulting nodes [33], malicious behavior from certain FL nodes [52, 53], and instances of the GAN attack in FL [54]. Other positive effects reported include improved accuracy in [55, 36, 56] and enhanced learning efficiency [57]. It must be noted that such benefits stem from incorporating XAI as a component of the FL algorithm design.
With the aim of detecting malicious attacks on FL operations, [53] uses a random forest (RF) to identify features causing incorrect predictions. Each participant trains both a DL and an RF model on their training data. For samples that are misclassified, it calculates the feature importance of each feature regarding incorrect classification, from all decision trees and using LIME [49]. The change in a feature’s importance over time is used to assign the contribution of each feature in misclassifying the data. Such an importance value provides insights into the level of influence of each log key in a sequence during an attack, thereby indicating which features should be most protected.
A novel FL protocol was proposed in [36] where a subset of all centers available online were selected to participate in each FL round, using Shapley values (computed using SHAP) as a heuristic to estimate FL contributions from each client. More specifically, at each round, the difference between local (feature-wise) and global aggregated Shapley values at each node is used to select participating parties in an FL process. The paper reports that the proposed method improves both the efficiency and accuracy compared to the FedAvg protocol [10], which does not consider client contributions.
The idea in [52] is to use explanation methods (specifically Backpropagation, GuidedBP, DeepLift, GradCAM [39], and Integrated Gradients) to detect whether each participant is using malicious data to enact a backdoor attack. To this end, so-called detection filters are developed. These consist of a classifier and an explanation method that respectively identify a likely backdoor attack and triggering features in the input data. The effectiveness of various explanation methods with different classifiers is tested, strengthening the FL process against backdoor attacks. Upon testing the detection accuracy in the presence of variable proportions of backdoor attacks, the proposed methodology proves the usefulness of the different explanation methods for backdoor attack prevention.
Adaptive sparse deep networks are implemented in [57], where parameters are shared via a multi-level federated network. At each round, weights are shared at the “top” and “second” sharing levels of the FL architecture, depending on the relevance values of the network calculated through Layerwise Relevance Propagation (LRP) [58]. This approach provides good diagnostic results even when the FL dataset is under a non-independent identical distribution (NOIID).
IV-C FL impacting model interpretability and vice versa
This subsection examines the relationship between Federated Learning (FL) and model interpretability. The featured studies, [59] and [60], discuss both the benefits and challenges of interpretability in distributed learning environments. These examples illustrate the trade-offs between model complexity and transparency that are inherent in FL setups.
An aggregation of client-based attention weights is investigated in [59] for a threat-detection task in a cloud scenario. Using system logs as input data, each client predicts cyberattacks and computes local attention weights, which are claimed to enhance interpretability. The central server subsequently aggregates these attention weights to build a saliency map that provides insights into the impact of the different log keys on the threat prediction. The influence of FL on model interpretability is assessed by comparing insights and interpretations from attention mechanisms for individual cyberattacks per client, individual attacks per aggregated models, and aggregated attacks per aggregated model. The aggregated insights proved to be more general, but at the expense of reduced interpretability.
In [60], a fuzzy rule-based system (FRBS) [61] is trained in federation via a one-shot communication scheme. Each data silo computes their own FRBS, and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule so the inferred regression function is piecewise linear, which is inherently explainable. It is observed in this work that FL and interpretability impact each other. The authors report that the average number of model rules is higher when trained in the federated setting (vs. centralized). Following the intuitive idea that higher complexity implies lower interpretability, this means that training via FL implies a lower interpretability in this specific study. On the other hand, the inherent interpretability of the models used in this study’s approach (Takagi-Sugeno-Kang FRBS [62]) allows the central server to identify conflicts between rules inferred by data centers. This approach facilitates the reconciliation of such discrepancies, thereby enhancing the learning process. This reconciliation not only improves model performance but also results in a more interpretable model compared to the ensemble of local models.
V Results
In this section, the information extracted from the papers included in the review is presented in a graphical form, to provide an overview of the different fields of application, explainability techniques, FL setups, and data types and modalities.
Figure 3: Field of application.
Fig. 3 shows the proportion of papers across the different fields of application, where medicine and life sciences are predominant (40.5%), followed by finance (16.2%), cybersecurity (13.5%), telecommunication (8.1%) and optimization in other engineering fields (8.1%) such as electricity, mechanics, and transportation. A small proportion (13.5%) of the analyzed papers focused on the theoretical side of the ML techniques without mentioning any specific application.

### v2

@@ Abstract
The joint implementation of federated learning (FL) and explainable artificial intelligence (XAI) could allow training models from distributed data and explaining their inner workings while preserving essential aspects of privacy. Toward establishing the benefits and tensions associated with their interplay, this scoping review maps the publications that jointly deal with FL and XAI, focusing on publications that reported an interplay between FL and model interpretability or post-hoc explanations. Out of the 37 studies meeting our criteria, only one explicitly and quantitatively analyzed the influence of FL on model explanations, revealing a significant research gap. The aggregation of interpretability metrics across FL nodes created generalized global insights at the expense of node-specific patterns being diluted. Several studies proposed FL algorithms incorporating explanation methods to safeguard the learning process against defaulting or malicious nodes. Studies using established FL libraries or following reporting guidelines are a minority. More quantitative research and structured, transparent practices are needed to fully understand their mutual impact and under which conditions it happens.
@@
Artificial Intelligence (AI), Data Privacy Preservation, Explainable AI, Federated Learning, Machine Learning, Model Interpretability.
@@ I Introduction
The development of trustworthy AI systems depends on different ethical principles like privacy preservation and explicability [1, 2]. Upholding these principles is essential to foster trust, ensure compliance, and maintain ethical integrity, particularly in data-sensitive fields like banking [3] and healthcare [4, 5].
Privacy concerns often limit AI models in healthcare to data from single institutions. Anonymization of large-scale healthcare data is difficult due to the risk of patient re-identification [6]. Federated learning (FL) helps solve this problem by allowing model training across multiple institutions without sharing raw data [7, 8]. FL enables machine learning (ML) models to train on distributed data while protecting privacy. It also addresses governance concerns by restricting direct access to sensitive information [9, 10]. Training on data from multiple institutions improves model generalizability and reduces bias compared to models trained on homogeneous datasets [11].
FL is categorized into three main types: horizontal FL (HFL), vertical FL (VFL), and federated transfer learning (FTL). HFL involves multiple clients holding different subsets of data points with the same feature space. VFL applies when institutions share data samples but with different feature spaces. FTL facilitates knowledge transfer across different FL settings, adapting models to new environments [12].
Explicability is another key ethical principle emphasized by ethicists and policymakers in high-risk AI domains like healthcare [2]. It requires that (a) AI processes be transparent, (b) the capabilities and purpose of AI systems be clearly communicated, and (c) AI decisions and predictions be explainable to those directly or indirectly affected [1, 2]. AI models that are transparent and understandable are more likely to earn the trust and acceptance of stakeholders, including customers, regulators, and users [13]. According to European Union regulations, aspects (a) and (c) of explicability, technically referred to as explainability [14], must be supported by additional information on how an AI system arrives at its outputs beyond performance metrics alone [15]. Explainable artificial intelligence (XAI) focuses on developing methods that make ML models more understandable. This can be achieved through interpretable model architectures or post-hoc explanation methods, aligning with requirements (a) and (c) of the EU guidelines [16].
However, the XAI literature lacks a consensus on the definition of explainability. Terms like explainability, explicability, and interpretability are often used interchangeably [17]. Some distinctions exist between interpretable modeling and explanation methods [15], which we adopt here due to their growing importance in high-stakes AI applications like healthcare [18]. An ML model is considered interpretable if its decision process can be inherently and intuitively understood by the intended user [15]. The term inherently is crucial, as interpretability is a passive property that reflects how naturally a model’s decisions make sense to a human observer [19] without requiring additional computation. Examples include linear and logistic regression models, where variable importance is inferred from model weights, and decision trees, which humans can intuitively follow [15]. In contrast, explanation methods actively perform additional computations on non-interpretable models to clarify their internal functions [19, 15]. These computations may involve the evaluation of model gradients for perturbed inputs [20], assessing feature importance [21], analyzing how output varies with feature changes [22], or identifying minimal input changes that alter predictions [23]. Explanation methods serve as interfaces between humans and AI systems while approximating the decision-making process [24].
Extensive research has been conducted on FL and XAI separately [25, 26], but their combined role in trustworthy AI remains underexplored. It is unclear whether FL and XAI complement each other, impose conflicting requirements, or can be addressed independently. XAI can help mitigate risks in federated learning by addressing challenges associated with decentralized data. Learning from heterogeneous datasets can introduce spurious correlations, as individual subsets may contain biases that distort the representation of the overall data distribution. Explainability helps practitioners assess whether model predictions are based on valid patterns rather than misleading correlations, reducing unintended biases and ensuring ethical compliance. Additionally, explainability can aid in detecting malicious agents attempting to poison the FL process. However, the interplay between FL and XAI entails potential risks that should be thoroughly examined and documented. There is no unified effort to determine whether FL reduces interpretability or affects the accuracy of explanations. Furthermore, explanations could expose vulnerabilities in the FL network, increasing susceptibility to attacks. Preliminary research [27] suggests that model explanations differ between FL and traditional centralized models, highlighting unique challenges in applying XAI in FL contexts. While many studies highlight the benefits of both perspectives, few quantitatively analyze their mutual impact. The literature also lacks a systematic examination of methodologies, challenges, and outcomes related to the interplay of FL and XAI, leaving a gap in understanding their interaction.
@@
To understand the interplay between FL and XAI, it is crucial to consider the findings, setups, and conditions under which these studies have been conducted. This scoping review maps experimental as well as methodological studies that explore their interaction. Since FL can potentially influence model interpretability and post-hoc explanations, the review investigates both aspects of XAI. We focus on research articles that either propose FL methods incorporating explainability, discuss practical applications where FL and XAI are jointly used, or analyze the interplay between these technologies through empirical studies. Given the fragmented and emerging nature of research at this intersection, this review provides a comprehensive overview of current work, identifies key concepts and evidence types, and highlights gaps to guide future research efforts.
@@
The remainder of this paper is structured as follows. Section II introduces the related work and Sec. III describes the method we followed. Analysis of the information extracted from the selected studies is presented in Sec. IV, and a more detailed analysis of the interplay between FL and XAI found in the reviewed papers is provided in Sec. IV-A. Section V discusses the implications of the reported results before Sec. VI concludes the paper. The Appendix provides a summary of how all included studies jointly address FL and XAI.
@@ II Related Work
Review papers focusing separately on XAI and FL are abundant, and the literature even includes several meta-reviews [25] and systematic reviews from different points of view (see, e.g., [28, 29] regarding FL applications for biomedical data). An extensive review of XAI methods can be found in [16] or [30]. Although there are a few reviews concerning the joint application of XAI and FL, none analyzes the interplay in as much depth as outlined in the following sections.
@@ II-A Reviews about FL and XAI
The review in [31] introduces Fed-XAI, which involves both the learning of interpretable models within a federated setting and the application of explanation methods to any federated model. They use [19] as a source for XAI taxonomy and highlight the diversity of definitions used in the studied literature. Additionally, they give an overview of the “current status in Fed-XAI” by discussing a set of articles relating to the two concepts with no analysis of the interplay between FL and XAI.
The review in [32] gathers relevant articles concerning the “FED-XAI” concept, defined by them as a discipline “which aims to bring together these two approaches into one”. Differing from our search terms, which include variations of the words “explanation” and “interpretability”, their search is limited to the explicit string “FED-XAI”, returning a reduced number of publications. The manuscript lacks a comparative analysis of the interplay between FL and XAI among the found articles.
The review in [33] surveys “interpretable federated learning” and proposes an interpretable FL taxonomy that enables learning models to explain prediction results, support model debugging, and provide insights into data owner contributions. The survey analyzes representative interpretable FL approaches, commonly adopted performance evaluation metrics, and future research directions. The primary focus is on leveraging interpretable methods to improve the FL process, but not on the impact of FL on the explanations of the predictions.
In contrast to these works, our review 1) encompasses an ample set of articles through a scoping review methodology, 2) provides detailed results about how FL and XAI are applied in each study, 3) summarizes how the reviewed studies jointly deal with FL and XAI, and 4) analyzes the findings on different kinds of interplay between them.
II-B Contribution-Aware Federated Learning
A subset of the approaches investigating the intersection of FL and XAI has utilized feature relevance methods (e.g., Shapley values) to measure the contribution of each client participating in an FL process. These approaches, commonly referred to as contribution-aware FL, aim to incentivize clients to engage during model training and to distribute rewards fairly based on contributions. One of the first approaches to contribution-aware FL suggested using Shapley values to interpret contributions in FL networks in [34]. The contributions are diverse, ranging from node liability [35] to biases within data sets [36]. Recent approaches have also extended SHAP-based [21] contribution determination to provide visualizations to evaluate data privacy [37] and to improve prediction reliability via clustering patients [38]. Such methods do not necessarily constitute an influence of XAI on the FL process and are, therefore, not central to the present review.
@@ III Methods
The present scoping review addresses the joint application of FL and XAI and their potential interplay. We aimed to examine the range and nature of research activity on the topic, summarize the research findings, and identify research gaps, following a standard scoping review methodology [39]. The presence of two concepts of explainability (interpretable models and post-hoc explanations), the two research types (methodological and experimental), and the diversity of joint approaches to FL and XAI increased the complexity of the study. Understanding and summarizing the diverse approaches required a deeper analysis of each included article. A more detailed description of the methodology is available in the pre-published research protocol [40].
III-A Identifying Research Questions
The research questions of this work address the training of interpretable ML models using FL and the explanation methods applied to ML models trained via FL. Furthermore, we categorized the research questions into two groups:
@@ III-A1 Methodological Advances
  • Has any study proposed an FL method that takes into account the ulterior application of explanation methods?
@@ III-A2 Experimental Results
  • In which contexts (fields of application) have FL and XAI been jointly evaluated?
@@
  • Has any study compared the effects of FL on the interpretability of the resulting models against a centralized learning algorithm?
  • Has any study compared the explanations obtained from a model trained via FL against a centralized learning algorithm?
III-B Identifying Relevant Studies
We tried to gather a broad range of manuscripts. Therefore, we used the five databases IEEEXplore, Google Scholar, PubMed, Scopus, and Web of Science. The terminological ambiguity between explainability and interpretability influenced the search strategy, which aimed to identify as many articles as possible that mention FL in conjunction with the (broad) concept of explainability. The search terms are detailed in Table I.
@@ III-B Identifying Relevant Studies
IEEEXplore (“federated learning”) AND ((explainable) OR (interpretable) OR (explaining) OR (explainability) OR (interpreting) OR (interpretability) OR (interpret))
@@
PubMed (“federated learning”) AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)
Scopus TITLE-ABS({federated learning} AND (explainable OR explaining OR explainability OR interpret OR interpreting OR interpretable OR interpretability)) AND (LIMIT-TO(DOCTYPE, “ar”) OR LIMIT-TO(DOCTYPE, “cp”))
@@
III-C Study Selection
We initially identified research articles published in peer-reviewed conferences and journals and works made available in pre-print services such as arXiv over the past three years. We excluded theses, reviews, meta-reviews, and surveys since they aim to summarize existing efforts and are discussed above. In our survey, the time frame of the published articles was between 2019 and April 2023 (date of protocol pre-publication). No relevant studies were found before 2019. We required that the full text be available. A flowchart summarizing the number of articles found, excluded, and included in this study, along the lines of [41], is displayed in Fig. 2.
To be included in the review, research articles had to discuss the relation and interaction between XAI and FL or apply both technologies jointly in a practical setting. We performed a screening procedure to rule out papers that only mention both technologies without detailed discussion or practical application. To that end, articles must answer positively to at least one of the following screening questions:
  • SQ1) Does the article propose a novel method integrating FL and XAI?
  • SQ2) Does the article report results from experiments applying FL and XAI in a real dataset?
  • SQ3) Does the article discuss or assess the impact of FL on explanations or model interpretability?
III-D Charting the Data
We utilized a shared online spreadsheet to facilitate data extraction. During both the screening and extraction process, double coding was practiced with the involvement of all authors. Each screened paper was independently assessed by two authors, and the extraction points for each included study were likewise completed independently by two authors. We tried resolving differences in data extraction through discussion between the two involved authors. If the disagreements persisted, a discussion within the entire author consortium resulted in the final decision. While extracting, we distinguished between explanation methods and model interpretability according to the nomenclature introduced above, even though some articles did not agree with the latter. We collected the following data:
@@ III-D Charting the Data
  • Nomenclature: whether the article provides or cites a definition of explainability, and whether the nomenclature regarding explainability coincides with the one introduced here.
  • Nature of the study: whether it is general, theoretical, or applied; if applied, what field it is applied on, and whether the proposed idea is practically validated in the field of application.
  • Data characteristics: information modality, type, and amount (number of unique data samples) of the data used in the experiments.
@@
  • XAI-specific characteristics: whether an interpretable model or an explanation method is used [^1] If the nomenclature in an article does not agree with the one mentioned in the Introduction, we identify whether it uses explanation methods or interpretable models according to our understanding., its type, and the specific explanation method (e.g.: SHAP, GradCAM [42]) or way of interpreting the model (e.g., weights at the first layer of DNN).
  • Interplay between FL and XAI: a paragraph summarizing how the article deals jointly with FL and XAI; influence of FL on explanations or interpretability (e.g. significant differences in variable importance ratings between centrally learned and model trained via FL); influence of XAI onto federated training (e.g., modified merging step in the central server after collecting locally trained models); and whether this influence is quantified, and how.
@@
IV Results
In this section, the information extracted from the 37 articles included in the review is collated, summarized, and reported. First, we give an overview of the studies where FL and XAI are applied together before presenting a more detailed description of the 11 articles where an interplay between FL and XAI was reported. Summaries of how the articles selected for review jointly deal with FL and XAI are provided in Appendix A.